OpenStudy (anonymous):

Solve the equation. 2x^2+7x=4 Give 2 answers for x.

OpenStudy (anonymous):

First step...get everything on the left side and 0 on the right. 2x^2 + 7x = 4 -4 -4 2x^2 + 7x -4 = 0 Next step...factor the left side. Do you know how to do that?

OpenStudy (anonymous):